OpenStudy (albert0898):

HELP 112. If k is a positive integer, then for every value of k, the sum of the k smallest distinct odd positive integers is equal to which of the following? F. k G. 2k H. 2k - 1 J. k^2 K. 2k^2 JUSTIFY YOUR ANSWER, HELP ME UNDERSTAND! Thank you!

OpenStudy (abb0t):

You should try taking your own tests next time.

OpenStudy (albert0898):

@abb0t Not a test. Actually studying for one.

OpenStudy (anonymous):

if k = 3, then 1 + 3 + 5 if k = 4, then 1 + 3 + 5 + 7 if k = 5, then 1 + 3 + 5 + 7 + 9 if k = 6, then 1 + 3 + 5 + 7 + 9 + 11 can you guess what the sum is?